A company needs to borrow $200,000 for 6 years. Barring any

cash flow constraints (i.e. the firm might find it easier to budget

monthly payments as opposed to semi annual payments), which

of the following two possible sources is the best option for the

firm?

1.

One source will lend them money at a rate of 16% compounded

monthly if the loan is paid by monthly payments.

2.

A second source will lend the money with semi-annual payments at

a rate of 18% compounded annually.
Answer #1

1. P=$200,000

t= 6 years

rate= 16% compounded monthly

A= P(1+r/n)^(nt)

A= 200000(1+0.16/12)^(12*6)= $519,036.23

total interest payable= $519,036.23-$200,000=$319,036.23

2. P= $200,000

t= 6 years

rate = 18% compounded annually

  A= P(1+r/n)^(nt)

A=200000(1+0.18)^6 = $539,910.83

interest payable= $539,910.83-$200,000=$339,910.83

the option 1 is best for the firm as it's payable interest is much less than the 2nd option,

the monthly payment is much less than the semi annual payment of 2nd option.